# Lense-Thirring precession of accretion flows: relativistic versus classical precession

### v1

@@
Lense-Thirring precession of accretion flows:
@@
Gabriel Török[Uncaptioned image] 11    Martin Urbanec[Uncaptioned image] 11    Monika Matuszková[Uncaptioned image] 11    Gabriela Urbancová[Uncaptioned image] 11    Kateřina Klimovičová[Uncaptioned image] 11    Debora Lančová[Uncaptioned image] 1122    Eva Šrámková[Uncaptioned image] 11    Jiří Horák[Uncaptioned image] 33
The vertical (Lense-Thirring) precession of the innermost accretion flows has been discussed as a sensitive indicator of the rotational properties of neutron stars (NSs) and their equation of state because it vanishes for a non-rotating star. In this work, we apply the Hartle-Thorne spacetimes to study the frequencies of the precession for both geodesic and non-geodesic (fluid) flows. We build on previous findings on the effect of the NS quadrupole moment, which revealed the importance of the interplay between the relativistic and classical precession. Because of this interplay, the widely used Lense-Thirring metric, linear in the NS angular momentum, is insufficient to calculate the behaviour of the precession frequency across an astrophysically relevant range of NS angular momentum values. We find that even for a moderately oblate NSs, the dependencies of the precession frequency on the NS angular momentum at radii within the innermost accretion region have maxima that occur at relatively low values of the NS angular momentum. We conclude that very different groups of accreting NSs, slow and fast rotators, can display the same precession frequencies. This may explain the lack of evidence for a correlation between the frequencies of the observed low-frequency quasiperiodic oscillations and the NS spin. In our work, we provide a full, general description of precession behaviour, and also examples that assume specific NS and quark star (MIT bag) equation of state. Our calculations are reproducible using the associated Wolfram Mathematica notebook.
@@ 1 Introduction
Neutron star low-mass X-ray binaries (NS LMXBs), which consist of an NS accreting matter from a less massive companion star, are invaluable for studying extreme environments and fundamental physics of dense matter (van der Klis, 2006; Lewin & van der Klis, 2010). Among the various phenomena observed in these systems, the X-ray variability exhibits a rich and diverse phenomenology.
@@
On short timescales, NS LMXBs often display a complex of quasi-periodic oscillations (QPOs) in the low- and high-frequency (LF and HF) bands. These features, represented by multiple peaks in the Fourier power-density spectra, span from units to a few tens of Hz for LF QPOs and from a few hundred Hz up to 1.3 kHz for HF QPOs. The HF QPOs often consist of two peaks, one of which is sometimes highly coherent, with quality factors (the height of the peak at its half-width) reaching up to $Q\sim 300$, and high root-mean-squared amplitude reaching up to $30\%$ (Barret et al., 2005b, 2006; Barret & Vaughan, 2012). The peaks within the LF QPO complex have relatively low coherence but still notable amplitudes (Barret et al., 2005a; Méndez, 2006; Barret & Vaughan, 2012). Similar behaviour is observed in the black hole LMXBs. However, the QPOs are weaker there, and the HF QPOs represent somewhat elusive features with small amplitudes and coherences. Commonly to both classes of sources, HF and LF QPO frequencies seem to be well correlated (Stella & Vietri, 1999; Mauche, 2002; van der Klis, 2006; Remillard & McClintock, 2006; Belloni & Motta, 2016).
@@
Most frequently, it is assumed that the QPOs emerge from the orbital motion in the vicinity of the central NS (Barret et al., 2006; van der Klis, 2006). A large variety of orbital models of QPOs have been proposed assuming either local effects within the accretion disc or the global motion of the disc itself (e.g., Kluźniak et al., 1990; Karas, 1999; Čadež et al., 2008; Kato, 2001; Wagoner et al., 2001). Many of them deal with an inner thick accretion flow (torus) and its epicyclic oscillations (Abramowicz & Kluźniak, 2001; Kluźniak & Abramowicz, 2001), or other modes of torus oscillations (Rezzolla et al., 2003a, b; Blaes et al., 2006; Fragile et al., 2016; Ingram & Done, 2010; Mishra et al., 2017; de Avellar et al., 2018).
@@
At the end of the millennium, the relativistic precession (RP) model was proposed (Stella & Vietri, 1998, 1999; Stella et al., 1999), which relates the HF and LF QPOs to the radial (periastron) and vertical (Lense-Thirring, LT) precession of inhomogeneities orbiting within the innermost accretion flow. They noticed that the vertical precession emerge as a critical tool for probing the characteristics and internal structure of NSs. Notably, the precession driven by the relativistic frame-dragging effects of the NS’s rotation should be absent in non-rotating NSs, making it possibly a distinctive indicator of NS rotation and equation of state (EoS). The concept of the RP model was later discussed in many works, including Ingram et al. (2009); Ingram & Done (2010, 2011); Middleton et al. (2019); Fragile (2020); Bollimpalli et al. (2024), where the authors focused on the vertical precession along with related results of numerical simulations of oscillations of accretion flows and their comparison with observations of LF QPOs.
@@
The predictions of the RP model and models predicting similar observable frequencies have been extensively studied in relation to both HF and LF QPOs in a variety of astrophysical sources (Stella & Vietri, 1998, 1999; Motta et al., 2012, 2014, 2022; Stuchl´ık & Kološ, 2016a, b; Török et al., 2016a, b, 2018; van Doesburgh et al., 2018; van Doesburgh & van der Klis, 2020; Kotrlová et al., 2020; Maselli et al., 2020). While in the former case, the expected frequencies in NS sources seem to agree with the observations (Belloni et al., 2007; Lin et al., 2011; Török et al., 2022), in the latter case, a lack of strong correlation between the frequencies of the LF QPOs and the NS spin was found (Altamirano et al., 2012; van Doesburgh & van der Klis, 2017, 2019; du Buisson et al., 2021).
@@ 2 Approach to rotating oblate compact stars
In the context of the RP model, the vertical epicyclic frequency and the frequency of LT precession of the geodesic motion around rotating oblate NSs were investigated by Morsink & Stella (1999), who calculated the associated post-Newtonian formulae, studied numerical models of rotating NSs, and compared the model predictions with the data. The behaviour of the precession frequencies and possible observable signatures of the NS structure were later investigated by Pappas (2012), who focused on the effects of multipole moments of spacetime. The study emphasised that, even for realistic NS models, the dependence of the precession frequency on the rotational frequency of the NS can be non-monotonic, and that the vertical epicyclic frequency can be greater than the Keplerian (see also Pappas & Apostolatos, 2012; Pappas, 2015).
The complex physics behind of the behaviour of the geodesic epicyclic frequencies around rotating oblate compact stars was further explored by the works of Kluźniak & Rosińska (2013, 2014), and Gondek-Rosińska et al. (2014), who , motivated by works of Zdunik & Gourgoulhon (2001) and Amsterdamski et al. (2002), calculated analytical formulae describing the epicyclic frequencies of rotating oblate bodies in Newtonian gravity, namely the Maclaurin spheroids, and compared them with those calculated for rotating quark stars (QS) using numerical general-relativistic approach. They showed the importance of the interplay between the relativistic effects and the effects arising purely from the spatial distribution of the gravitating mass. In terms of the LT precession, this interplay manifests itself as a competition between the relativistic and classical precession. The results of these works were consequently extended in some aspects by Tsang & Pappas (2016), who focused on the self-trapping of discoseismic corrugation modes in NS spacetimes, and Bollimpalli (2022), who calculated epicyclic frequencies in the case of Newtonian stars composed of concentric spheroids.
In our study, we generalise the previous results and study the LT precession in a fully analytic general-relativistic approach that assumes a rotating oblate NS and frequencies associated with both geodesic motion and non-geodesic fluid flows.
@@ 2.1 Spacetime
We study space-time geometry around a rotating, axisymmetric compact object governed by general relativity in terms of the Schwarzschild coordinates $\{t,r,\theta,\varphi\}$. The corresponding metric tensor captures the curvature induced by the star’s mass and rotation, and is represented by [^1] In the adopted framework, the units of measurement are chosen so that both the speed of light $c$ and the gravitational constant $G$ are equal to one ($c=G=1$). Distances are measured in terms of the gravitational radius, $r_{\mathrm{G}}=GM/c^{2}$, and the metric signature used is (,+,+,+)(-,+,+,+).:
@@
To handle the calculation on the NS space-time background, the LT metric Lense & Thirring (1918) is often used. However, this metric is only linear in the NS angular momentum. Here we adopt the Hartle-Thorne metric (HT, Hartle, 1967; Hartle & Thorne, 1968), which takes into account the star’s mass $M$, angular momentum $J$, and quadrupole moment $Q$. To simplify the expressions, we use the dimensionless angular momentum and quadrupole moment, $j=J/M^{2}$ and $q=Q/M^{3}$. The components of the HT metric in notation adopted here (Abramowicz et al., 2003; Urbancová et al., 2019) are, along with the components of the linear metric, given in a Wolfram Mathematica notebook associated with this work. [^2] https://github.com/Astrocomp/LTprecessionHT
@@ 2.2 Quadrupole parameter
In principle, a non-rotating NS can have non-zero oblateness. Here, we assume the oblateness to be entirely rotationally induced and thus to vanish when the angular momentum of the star reaches zero. To study the effects of the oblateness and angular momentum on the orbital motion, we utilise the quadrupole parameter,
@@
This quantity is determined by the given NS EoS and $M$ and does not depend on the NS rotational frequency (e.g., Urbancová et al., 2019). Furthermore, in terms of the space-time geometry, its value determines whether the influence of the $g_{t\varphi}$ component of the metric tensor and the associated frame-dragging effects, which are linear in $j$, dominate or can be overtaken by the other components. This occurs as $j$ increases, and as the effects given purely by the spatial distribution of the gravitating mass associated with $j^{2}$ become important.
2.3 Range of Space-Time Parameters
The range of parameters considered here is constrained by the observed properties of neutron stars in accreting systems, with masses ranging from $1.4M_{\odot}$ to $2.5M_{\odot}$ (Strohmayer et al., 1996; Lattimer & Prakash, 2001) and rotational frequencies corresponding to millisecond pulsars. Previous studies of NS models have shown that the maximum dimensionless angular momentum is approximately $j_{\text{max}}\approx 0.7$, with a corresponding quadrupole parameter $\tilde{q}$ ranging from $\tilde{q}\leavevmode\nobreak\ \sim\leavevmode\nobreak\ 1.5$ for a very massive (compact) NS to $\tilde{q}\leavevmode\nobreak\ \sim\leavevmode\nobreak\ 10$ for a low-mass NS depending on the NS equation of state (Lo & Lin, 2011; Urbancová et al., 2019).
@@ 3 Orbital motion
Figure 1: Frequencies of orbital motion around rotating NS. a) NS spacetime approximated using the LT metric. b) Spacetime of extremely compact NS approximated using the Kerr limit of the HT metric. c) Spacetime of highly oblate NS approximated by the HT metric. The upper panels show the Keplerian and (axisymmetric) epicyclic frequencies. The bottom panels show the precession frequencies. The full lines correspond to the frequencies of a test particle motion and slender tori. The colour-shaded regions indicate frequencies relevant to the non-geodesic flows of non-slender tori. Dashed lines in panels b indicate the frequencies of a test particle motion in the Kerr spacetime calculated using the appropriate coordinate transformation.
Figure 2: Change of the orientation of the LT precession and behaviour of local frequency maxima shown for $j=0.35$. As in Figure 1, the solid lines correspond to the frequencies of test particle motion and slender tori, while the colour-shaded regions indicate the frequencies relevant to non-geodesic flows. Full black circles indicate local maxima of the geodesic precession frequency, $\nu_{\mathrm{max}}^{0}$.
@@ 4.1 Fluid flows
To inspect the behaviour of the fluid flow oscillations, we follow Matuszková et al. (2024a, b) and consider an inner accretion torus orbiting an NS. The adopted torus model assumes a fluid with constant specific angular momentum throughout the flow, with Keplerian value at the torus centre. A polytropic equation of state further describes the fluid. In contrast to the oscillation frequencies of a free test particle, the oscillations of such fluid torus involve additional corrections due to the internal pressure forces. The oscillation frequencies are given as
@@ 4.2 Inner parts of the flow
Then, there is the radius close to ISCO, $r_{3:1}$, where $\nu_{\theta}=3\nu_{r}$, and the radius $r_{2:1}$, where $\nu_{\theta}=2\nu_{r}$ and $\nu_{r}$ reaches a maximum. When the NS spacetime differ from the case of $j=0$, for the maximum of $\nu_{r}$ there is still $\nu_{\theta}\approx 2\nu_{r}$ (Török et al., 2008). Last, we mark the radius where $2\nu_{\theta}=3\nu_{r}$. In next, we inspect how $\nu_{\mathrm{LT}}$ changes with the NS spin along these four radii, which covers the innermost parts of an accretion flow.

### v2

# Lense-Thirring precession of neutron-star accretion flows: Relativistic versus classical precession

@@
Lense-Thirring precession of neutron-star accretion flows:
@@
Gabriel Török[Uncaptioned image]    Martin Urbanec[Uncaptioned image]    Monika Matuszková[Uncaptioned image]    Gabriela Urbancová[Uncaptioned image]    Kateřina Klimovičová[Uncaptioned image]    Debora Lančová[Uncaptioned image]    Eva Šrámková[Uncaptioned image]    Jiří Horák[Uncaptioned image]
The vertical (Lense-Thirring) precession of the innermost accretion flows has been discussed as a sensitive indicator of the rotational properties of neutron stars (NSs) and their equation of state because it vanishes for a non-rotating star. In this work, we apply the Hartle-Thorne spacetimes to study the frequencies of the precession for both geodesic and non-geodesic (fluid) flows. We build on previous findings on the effect of the NS quadrupole moment, which revealed the importance of the interplay between the relativistic and classical precession. Because of this interplay, the widely used Lense-Thirring metric, linear in the NS angular momentum, is insufficient to calculate the behaviour of the precession frequency across an astrophysically relevant range of NS angular momentum values. We find that even for a moderately oblate NSs, the dependencies of the precession frequency on the NS angular momentum at radii within the innermost accretion region have maxima that occur at relatively low values of the NS angular momentum. We conclude that very different groups of accreting NSs — slow and fast rotators — can display the same precession frequencies. This may explain the lack of evidence for a correlation between the frequencies of the observed low-frequency quasiperiodic oscillations and the NS spin. In our work, we provide a full, general description of precession behaviour, and also examples that assume specific NS and quark star (MIT bag) equation of state. Our calculations are reproducible using the associated Wolfram Mathematica notebook.
@@ 1 Introduction
Neutron star low-mass X-ray binaries (NS LMXBs), which consist of an NS accreting matter from a less massive companion star, are invaluable for studying extreme environments and fundamental physics of dense matter (van der Klis, 2006; Lewin and van der Klis, 2010). Among the various phenomena observed in these systems, the X-ray variability exhibits a rich and diverse phenomenology.
@@
On short timescales, NS LMXBs often display a complex of quasi-periodic oscillations (QPOs) in the low- and high-frequency (LF and HF) bands. These features, represented by multiple peaks in the Fourier power-density spectra, span from a few hertz to a few tens of hertz for LF QPOs and from a few hundred hertz up to 1.3 kHz for HF QPOs. The HF QPOs often consist of two peaks, one of which is sometimes highly coherent, with quality factors (the height of the peak at its half-width) reaching up to $Q\sim 300$, and high root mean square amplitude reaching up to $30\%$ (Barret et al., 2005b, 2006; Barret and Vaughan, 2012). The peaks within the LF QPO complex have relatively low coherence but still notable amplitudes (Barret et al., 2005a; Méndez, 2006; Barret and Vaughan, 2012). Similar behaviour is observed in the black hole LMXBs. However, the QPOs are weaker there, and the HF QPOs represent somewhat elusive features with small amplitudes and coherences. Commonly to both classes of sources, HF and LF QPO frequencies seem to be well correlated (Stella and Vietri, 1999; Mauche, 2002; van der Klis, 2006; Remillard and McClintock, 2006; Belloni and Motta, 2016).
@@
It is frequently assumed that the QPOs emerge from the orbital motion in the vicinity of the central NS (Barret et al., 2006; van der Klis, 2006). A large variety of orbital models of QPOs have been proposed assuming either local effects within the accretion disc or the global motion of the disc itself (e.g., Kluźniak et al., 1990; Karas, 1999; Čadež et al., 2008; Kato, 2001; Wagoner et al., 2001). Many of them deal with an inner thick accretion flow (torus) and its epicyclic oscillations (Abramowicz and Kluźniak, 2001; Kluźniak and Abramowicz, 2001), or other modes of torus oscillations (Rezzolla et al., 2003a, b; Blaes et al., 2006; Fragile et al., 2016; Ingram and Done, 2010; Mishra et al., 2017; de Avellar et al., 2018).
@@
At the end of the millennium, the relativistic precession (RP) model was proposed (Stella and Vietri, 1998, 1999; Stella et al., 1999), which relates the HF and LF QPOs to the radial (periastron) and vertical (Lense-Thirring, LT) precession of inhomogeneities orbiting within the innermost accretion flow. They noticed that the vertical precession emerges as a critical tool for probing the characteristics and internal structure of NSs. Notably, the precession driven by the relativistic frame-dragging effects of the NS’s rotation should be absent in non-rotating NSs, making it possibly a distinctive indicator of NS rotation and equation of state (EoS). The concept of the RP model was later discussed in many works, including Ingram et al. (2009); Ingram and Done (2010, 2011); Middleton et al. (2019); Fragile (2020), and Bollimpalli et al. (2024) where the authors focused on the vertical precession along with related results of numerical simulations of oscillations of accretion flows and their comparison with observations of LF QPOs.
@@
The predictions of the RP model and models predicting similar observable frequencies have been extensively studied in relation to both HF and LF QPOs in a variety of astrophysical sources (Stella and Vietri, 1998, 1999; Motta et al., 2012, 2014, 2022; Stuchlík and Kološ, 2016a, b; Török et al., 2016a, b, 2018; van Doesburgh et al., 2018; van Doesburgh and van der Klis, 2020; Kotrlová et al., 2020; Maselli et al., 2020). While in the former case the expected frequencies in NS sources seem to agree with the observations (Belloni et al., 2007; Lin et al., 2011; Török et al., 2022), in the latter case a lack of strong correlation between the frequencies of the LF QPOs and the NS spin was found (Altamirano et al., 2012; van Doesburgh and van der Klis, 2017, 2019; du Buisson et al., 2021).
@@ 2 Approach to rotating oblate compact stars
In the context of the RP model, the vertical epicyclic frequency and the frequency of LT precession of the geodesic motion around rotating oblate NSs were investigated by Morsink and Stella (1999), who calculated the associated post-Newtonian formulae, studied numerical models of rotating NSs, and compared the model predictions with the data. The behaviour of the precession frequencies and possible observable signatures of the NS structure were later investigated by Pappas (2012), who focused on the effects of multipole moments of spacetime. The study emphasised that, even for realistic NS models, the dependence of the precession frequency on the rotational frequency of the NS can be non-monotonic, and that the vertical epicyclic frequency can be greater than the Keplerian (see also Pappas and Apostolatos, 2012; Pappas, 2015).
The complex physics behind the behaviour of the geodesic epicyclic frequencies around rotating oblate compact stars was further explored in the works of Kluźniak and Rosińska (2013, 2014), and Gondek-Rosińska et al. (2014), who , motivated by the works of Zdunik and Gourgoulhon (2001) and Amsterdamski et al. (2002), calculated analytical formulae describing the epicyclic frequencies of rotating oblate bodies in Newtonian gravity, namely the Maclaurin spheroids, and compared them with those calculated for rotating quark stars (QSs) using a numerical general-relativistic approach. They showed the importance of the interplay between the relativistic effects and the effects arising purely from the spatial distribution of the gravitating mass. In terms of the LT precession, this interplay manifests itself as a competition between the relativistic and classical precession. The results of these works were consequently extended in some aspects by Tsang and Pappas (2016), who focused on the self-trapping of discoseismic corrugation modes in NS spacetimes, and Bollimpalli (2022), who calculated epicyclic frequencies in the case of Newtonian stars composed of concentric spheroids.
Here, we generalised the previous results. Using a fully analytic, general relativistic approach, we examined LT precession in relation to any rotating, oblate compact stars, including NSs and QSs. Furthermore, we investigated not only frequencies associated with geodesic motion, but also those associated with non-geodesic fluid flows.
@@ 2.1 Spacetime
We study spacetime geometry around a rotating, axisymmetric compact object governed by general relativity in terms of the Schwarzschild coordinates $\{t,r,\theta,\varphi\}$. The corresponding metric tensor captures the curvature induced by the star’s mass and rotation, and is represented by [^1] In the adopted framework, the units of measurement are chosen so that both the speed of light $c$ and the gravitational constant $G$ are equal to one ($c=G=1$). Distances are measured in terms of the gravitational radius, $r_{\mathrm{G}}=GM/c^{2}$, and the metric signature used is (,+,+,+)(-,+,+,+).:
@@
To handle the calculation on the NS spacetime background, the LT metric Lense and Thirring (1918) is often used. However, this metric is only linear in the NS angular momentum. Here we adopt the Hartle-Thorne metric (HT, Hartle, 1967; Hartle and Thorne, 1968), which takes into account the star’s mass $M$, angular momentum $J$, and quadrupole moment $Q$. To simplify the expressions, we use the dimensionless angular momentum and quadrupole moment, $j=J/M^{2}$ and $q=Q/M^{3}$. The components of the HT metric in notation adopted here (Abramowicz et al., 2003; Urbancová et al., 2019) are, along with the components of the linear metric, given in a Wolfram Mathematica notebook associated with this work. [^2] https://github.com/Astrocomp/LTprecessionHT
@@ 2.2 Quadrupole parameter
In principle, a non-rotating NS can have non-zero oblateness. Here, we assume the oblateness to be entirely rotationally induced, and thus to vanish when the angular momentum of the star reaches zero. To study the effects of the oblateness and angular momentum on the orbital motion, we utilise the quadrupole parameter,
@@
This quantity is determined by the given NS EoS and $M$ and does not depend on the NS rotational frequency (e.g., Urbancová et al., 2019). Furthermore, in terms of the spacetime geometry, its value determines whether the influence of the $g_{t\varphi}$ component of the metric tensor and the associated frame-dragging effects, which are linear in $j$, dominate or can be overtaken by the other components. This occurs as $j$ increases, and as the effects given purely by the spatial distribution of the gravitating mass associated with $j^{2}$ become important.
2.3 Range of spacetime parameters
The range of parameters considered here is constrained by the observed properties of neutron stars in accreting systems, with masses ranging from $1.4M_{\odot}$ to $2.5M_{\odot}$ (Strohmayer et al., 1996; Lattimer and Prakash, 2001) and rotational frequencies corresponding to millisecond pulsars. Previous studies of NS models have shown that the maximum dimensionless angular momentum is approximately $j_{\text{max}}\approx 0.7$, with a corresponding quadrupole parameter $\tilde{q}$ ranging from $\tilde{q}\penalty 10000\ \sim\penalty 10000\ 1.5$ for a very massive (compact) NS to $\tilde{q}\penalty 10000\ \sim\penalty 10000\ 10$ for a low-mass NS depending on the NS equation of state (Lo and Lin, 2011; Urbancová et al., 2019).
@@ 3 Orbital motion
Figure 1: Frequencies of orbital motion around rotating NSs. a) NS spacetime approximated using the LT metric. b) Spacetime of extremely compact NSs approximated using the Kerr limit of the HT metric. c) Spacetime of highly oblate NSs approximated by the HT metric. The upper panels show the Keplerian and (axisymmetric) epicyclic frequencies. The bottom panels show the precession frequencies. The full lines correspond to the frequencies of a test particle motion and slender tori. The colour-shaded regions indicate frequencies relevant to the non-geodesic flows of non-slender tori. The dashed lines in panel b indicate the frequencies of a test particle motion in the Kerr spacetime calculated using the appropriate coordinate transformation.
Figure 2: Change in the orientation of the LT precession and behaviour of local frequency maxima shown for $j=0.35$. As in Figure 1, the solid lines correspond to the frequencies of test particle motion and slender tori, while the colour-shaded regions indicate the frequencies relevant to non-geodesic flows. The full black circles indicate local maxima of the geodesic precession frequency, $\nu_{\mathrm{max}}^{0}$.
@@ 4.1 Fluid flows
To inspect the behaviour of the fluid flow oscillations, we follow Matuszková et al. (2024a, b) and consider an inner accretion torus orbiting an NS. The adopted torus model assumes a fluid with constant specific angular momentum throughout the flow, with Keplerian value at the torus centre. A polytropic equation of state further describes the fluid. In contrast to the oscillation frequencies of a free test particle, the oscillations of such a fluid torus involve additional corrections due to the internal pressure forces. The oscillation frequencies are given as
@@ 4.2 Inner parts of the flow
Then, there is the radius close to ISCO, $r_{3:1}$, where $\nu_{\theta}=3\nu_{r}$, and the radius $r_{2:1}$, where $\nu_{\theta}=2\nu_{r}$ and $\nu_{r}$ reaches a maximum. When the NS spacetime differ from the case of $j=0$, for the maximum of $\nu_{r}$ there is still $\nu_{\theta}\approx 2\nu_{r}$ (Török et al., 2008). Last, we mark the radius where $2\nu_{\theta}=3\nu_{r}$. Next, we inspect how $\nu_{\mathrm{LT}}$ changes with the NS spin along these four radii, which cover the innermost parts of an accretion flow.